# MetaCLBench: Meta Continual Learning Benchmark on Resource-Constrained Edge Devices

### v1

@@ 3.4 Model Architecture
YAMNet Howard et al. (2017) employs the MobileNetV1, which features a depthwise separable convolution architecture designed to decrease model size and latency and a full convolution on the first layer. The depthwise separable convolution comprises a deep convolution responsible for filtering and a pointwise convolution used $1\times 1$ convolution responsible for combining features, effectively mixing information from all output channels in the deep convolution step. MobileNetV1 consists of 28 layers, with each layer being downsampled through strided convolution. Additionally, MobileNets introduces two hyperparameters, a width multiplier and a resolution multiplier, which facilitate the balance latency and accuracy.
@@
ViT Dosovitskiy et al. (2020) embodies a hybrid architecture. Since ViT only has a Multilayer Perceptron (MLP) layer with localization and translation equivalence, the image-specific generalization bias of ViT is much smaller than that of CNN. The same five methodologies deployed in the 3-layer architecture have been adapted for use with ViT. ViT begins by partitioning an image into patches; these patches, extracted from the CNN feature maps, serve as alternative input sequences for the model.
3.5 System Implementation
@@
The implementation of our benchmark framework, MetaCLBench, consists of two stages. We developed the first stage, pre-training and meta-training, of Meta CL methods on a Linux server to initialize and optimize neural weights in order to enable fast adaptation during deployment scenarios with a few samples. After that, we implemented the second stage, meta-testing (i.e., actual deployment setup), on our target devices: (1) embedded and mobile systems such as Jetson Nano and Raspberry Pi 3B+ and (2) a severely resource-constrained IoT device like Raspberry Pi Zero 2. In addition, for methods that utilize hardware-friendly optimization (e.g., LifeLearner), we also adopt their optimization techniques in our framework. Specifically, MetaCLBench incorporates hardware-friendly 8-bit integer arithmetic Jacob et al. (2018) which reduces the precision of weights/activations of the model from 32-bit floats to 8-bit integers, increasing the computation throughput and minimizing latency and energy. MetaCLBench uses the scalar quantization scheme Jacob et al. (2018) to minimize the information loss in quantization. Also, the QNNPACK backend engine is used to execute the quantized model on the embedded devices.
Embedded Devices: MetaCLBench employs three diverse edge devices: (1) Jetson Nano, (2) Raspberry Pi 3B+, and (3) Raspberry Pi Zero 2 which are equipped with a quad-core ARM processor and 4 GB, 1 GB, 0.5 GB of RAM, respectively. In addition, it should be noted that the available memory on the Jetson Nano, Raspberry Pi 3B+, and Pi Zero 2 during idle time is about 1.7 GB, 600 MB, and 250 MB. This is because some of the memory is already being used by background processes, concurrent apps, and the operating system. Hence, we allocate an additional 1GB of swap space, enabling the execution of memory-intensive Meta-CL methods such as ANML+AIM and OML+AIM. Our software stack includes Faiss Johnson et al. (2019) and PyTorch 1.13 for meta-training and meta-testing stages.
3.6 Experimental Setup
@@
Training Detail: We adopt a pre-training procedure, similar to prior works Hu et al. (2022); Yosinski et al. (2014). For the advanced model architectures such as ViT and YAMNet, there exist pre-trained model weights, thus we utilize them directly similar to Hu et al. (2022). Then, for the CNN architecture, we pre-train the CNN model with a sufficient number of epochs (i.e., until the validation loss converges), which is consistent with conventional transfer learning for DNNs Yosinski et al. (2014). Consistent with prior meta-training research Lee et al. (2021); Kwon et al. (2024a), we employed a batch size of 1 for the inner-loop updates and 64 for the outer-loop updates across 20,000 steps to ensure the accuracy of our results. To obtain a meta-trained model with the highest possible validation set accuracy, we tested with various learning rates required for both the inner and outer loops. Consequently, we established the learning rate for the inner loop ($\alpha$) was set to 0.001, and the learning rate for the outer loop ($\beta$) was also set to 0.001 for all datasets. In the meta-testing phase, we assessed ten distinct learning rates for each method and reported the best accuracy. To evaluate the precision of rehearsal-based methods, we experimented with batch sizes of 8 and 16 and observed a minimal performance difference depending on the batch sizes. We selected a batch size of 8 as it requires less memory than larger batch sizes.
@@
Evaluation Metrics: Following Beaulieu et al. (2020), this study focuses on the accuracy of unseen samples across new categories in CL. We measure the end-to-end latency, energy consumption, and peak memory by deploying various Meta-CL methods using MetaCLBench on edge devices: Jetson Nano, Raspberry Pi 3B+, Raspberry Pi Zero 2. Specifically, we measure system performance metrics by continually learning all given classes for deployed DNNs on embedded devices. Peak memory includes: (1) model memory for updated weights, (2) optimizer memory for gradients and momentum, (3) activations memory for intermediate outputs during weight updates, and (4) rehearsal samples. To measure end-to-end latency and energy consumption, we account for the time and energy required to: (1) load the model and (2) perform Meta-CL using all available samples (e.g., 30) across all classes (e.g., 30 for CIFAR-100 and 10 for GSCv2). Energy consumption is measured by monitoring the power consumption of the edge device using a YOTINO USB power meter. We derive the energy consumption using the equation (Energy = Power x Time).
4 Results & Findings
@@
This section presents comprehensive evaluation results and findings from MetaCLBench. Due to page constraints, we present results from two datasets in the main text, with complete results available in the Appendix.
4.1 Main Results
@@
Table 1: Performance and computational costs, memory footprint of six representative Meta-CL methods using three network architectures on five datasets of image and audio domains. OOM indicates an out-of-memory issue.
@@
Pre-trained 0.258 474.5MB 1,198s 5.5KJ
@@
Oracle ANML 0.438 475.0 MB 3.414s 15.7KJ
LifeLearner 0.433 136.7 MB 1,236s 6.17kJ
Pre-trained 0.182 1,382 MB 302s 1.6KJ
@@
Urban ANML+AIM 0.439 2,593 MB OOM OOM
Sound8K OML+AIM 0.385 2,648 MB OOM OOM
Raw ANML 0.667 1,456 MB 60,120s 279KJ
@@
Pre-trained 0.234 336.6 MB 2,056s 11.2KJ
Mini OML 0.255 336.6 MB 2,203s 11.4KJ
ImageNet Raw OML 0.454 1,334 MB 301,456s 1150KJ
Oracle OML 0.512 337.7 MB 6,301s 31KJ
Latent OML 0.376 336.6 MB 2,250s 13KJ
@@
Pre-trained 0.186 39.2 MB 912s 5.4KJ
Urban OML 0.262 39.2 MB 910s 5.4KJ
Sound8K Raw OML 0.595 139.3 MB 120,965s 550KJ
Oracle OML 0.729 42.4 MB 2,501s 21.1KJ
@@
Previous research Kwon et al. (2024a); Holla et al. (2020), has shown that OML algorithms achieve lower accuracy than ANML algorithms when using 3-layer CNN architectures. Therefore, we focused our experimental investigation exclusively on ANML for CNN-based evaluations. For advanced models like YAMNet and ViT, we prioritized OML evaluation due to their inherent pre-training components, which eliminate the need for additional pre-training through ANML or the AIM module—additions that neither significantly improve accuracy nor justify their computational cost.
Accuracy. We evaluated the test accuracy of six Meta-CL approaches across five datasets against a baseline model. Oracle models (Oracle ANML for CNN, Oracle OML for YAMNet and ViT) establish the upper accuracy bounds and serve as benchmarks, as they are trained on all classes simultaneously in an i.i.d. manner, following traditional offline DNN training. By comparing Meta-CL methods against Oracle models, we demonstrate the trade-offs between Meta-CL approaches and ideal scenarios where all data is available initially, highlighting our methods’ efficiency under practical constraints. The results in Table 1 show that models using only pre-trained weights ("Pre-trained") achieve lower accuracy, averaging 22.18% and 13.8% across the datasets for Convolutional Neural Networks (CNNs) and advanced architectures(YAMNet and ViT). This result indicates that traditional transfer learning fails to address few-shot learning challenges. In CNN architecture, baseline ANML improved upon the pre-trained-only approach with a 16.7% average accuracy gain. The AIM enhancement further increased accuracy by 2.54%. However, these improvements remained below Oracle and LifeLearner methods by 15.92%. The constrained sample size during the meta-testing phase poses a challenge to attaining high accuracy, even considering the theoretical upper limit.
Peak Memory Footprint. We conducted measurements to ascertain the maximum RAM requirements forbackpropagation execution and rehearsal samples storage. The backpropagation memory is divided into three parts: (1) model memory, which is responsible for storing model parameters; (2) optimizer memory, which is responsible for storing gradient and momentum vectors; (3) activation memory, which is responsible for storing intermediate activations that are reused during backpropagation. Table 1 presents the peak memory usage across various methodologies. First, we found that although AIM enhances accuracy, it increases peak usage due to additional parameters and activation storage requirements during training. The memory demands 1,562 MB for MiniImageNet and 2,648 MB for UrbanSound8K, significantly exceeding the RAM capacity of embedded devices such as the Pi 3B+ and resulting out-of-memory (OOM) problems. Consequently, it becomes challenging to utilize it in edge devices. While OML, ANML, Oracle, Pre-trained, and LifeLearner demonstrate lower memory requirements, Pre-trained and ANML are less accurate. LifeLearner emerges as the optimal solution, requiring only 136-496 MB while maintaining accuracy comparable to Oracle through high compression rates. This efficiency is achieved through latent replay, which reduces activation memory and overall memory footprint, making it particularly suitable for resource-constrained environments. In contrast, advanced architectures like ViT and YAMNet demand greater computational resources due to high-dimensional data processing and additional layer training requirements.
@@
End-to-end Latency & Energy Consumption. Table 1 also presents the operational efficiency of Meta-CL methods implemented on Raspberry Pi 3B+, focusing on end-to-end latency and energy consumption. The end-to-end latency comprises four components: model loading (5%), inference (30%), backpropagation (61%), and data processing overhead (4%). First, we measured the end-to-end latency of CL methods on the Raspberry Pi 3B+ across all classes (30 samples per class). LifeLearner achieved fast end-to-end latency. Oracle exhibited higher training latency due to i.i.d. Training over multiple epochs until it converges. AIM’s implementation of attention mechanism selection during backpropagation increased computational overhead. LifeLearner reduces training and memory latency through latent replay and model freezing. Due to the large amount of memory required by the AIM series, crashes occurred several times during the test due to memory exhaustion.
To establish generalizability, we extended our evaluation to diverse edge devices with varying computational capabilities (Jetson Nano with enhanced specifications and Pi Zero 2 with reduced capacity). The table 2 below shows the end-to-end latency and energy consumption results across these three edge devices, confirming that LifeLearner achieves a 7.5-fold reduction in latency compared to Oracle.
@@
Table 2: End-to-end latency and energy consumption results of running different Meta-CL methods on three different edge devices

### v2

@@ III-D Model Architecture
YAMNet [13] employs the MobileNetV1, which features a depthwise separable convolution architecture designed to decrease model size and latency and a full convolution on the first layer. The depthwise separable convolution comprises a deep convolution responsible for filtering and a pointwise convolution used $1\times 1$ convolution responsible for combining features, effectively mixing information from all output channels in the deep convolution step. MobileNetV1 consists of 28 layers, with each layer being downsampled through strided convolution. Additionally, MobileNets introduces two hyperparameters, a width multiplier and a resolution multiplier, which facilitate the balance between latency and accuracy.
@@
ViT [9] embodies a hybrid architecture. Since ViT only has a Multilayer Perceptron (MLP) layer with localization and translation equivalence, the image-specific generalization bias of ViT is much smaller than that of CNN. The same five methodologies deployed in the 3-layer architecture have been adapted for use with ViT. ViT begins by partitioning an image into patches; these patches, extracted from the CNN feature maps, serve as alternative input sequences for the model.
III-E System Implementation
@@
The implementation of our benchmark framework, MetaCLBench, consists of two stages. We developed the first stage, pre-training and meta-training, of Meta CL methods on a Linux server to initialize and optimize neural weights in order to enable fast adaptation during deployment scenarios with a few samples. After that, we implemented the second stage, meta-testing (i.e., actual deployment setup), on our target devices: (1) embedded and mobile systems such as Jetson Nano and Raspberry Pi 3B+, and (2) a severely resource-constrained IoT device like Raspberry Pi Zero 2. In addition, for methods that utilize hardware-friendly optimization (e.g., LifeLearner), we also adopt their optimization techniques in our framework. Specifically, MetaCLBench incorporates hardware-friendly 8-bit integer arithmetic [16] which reduces the precision of weights/activations of the model from 32-bit floats to 8-bit integers, increasing the computation throughput and minimizing latency and energy. MetaCLBench uses the scalar quantization scheme [16] to minimize the information loss in quantization. Also, the QNNPACK backend engine is used to execute the quantized model on the embedded devices.
Target IoT Hardware: We evaluate on three device classes representing the IoT hardware spectrum:
  • Jetson Nano (4 GB): Entry-level edge AI platform, representative of smart cameras and local inference gateways.
  • Raspberry Pi 3B+ (1 GB): General-purpose IoT device, representative of smart home hubs and sensor gateways.
  • Raspberry Pi Zero 2 (512 MB): Severely constrained platform, representative of wearables and low-cost distributed sensors.
The gap between nominal and available RAM—caused by OS overhead and background services—is critical for deployment planning but often overlooked in ML benchmarks. Available memory during item time is approximately 1.7 GB, 600 MB, and 250 MB respectively. Hence, we allocate an additional 1GB of swap space, enabling the execution of memory-intensive Meta-CL methods such as ANML+AIM and OML+AIM. Our software stack includes Faiss [22] and PyTorch 1.13 for meta-training and meta-testing stages.
III-F Experimental Setup
@@
Training Detail: We adopt a pre-training procedure, similar to prior works [14, 56]. For the advanced model architectures such as ViT and YAMNet, there exist pre-trained model weights, thus we utilize them directly similar to [14]. Then, for the CNN architecture, we pre-train the CNN model with a sufficient number of epochs (i.e., until the validation loss converges), which is consistent with conventional transfer learning for DNNs [56]. Consistent with prior meta-training research [31, 26], we employed a batch size of 1 for the inner-loop updates and 64 for the outer-loop updates across 20,000 steps to ensure the accuracy of our results. To obtain a meta-trained model with the highest possible validation set accuracy, we tested with various learning rates required for both the inner and outer loops. Consequently, we established the learning rate for the inner loop ($\alpha$) was set to 0.001, and the learning rate for the outer loop ($\beta$) was also set to 0.001 for all datasets. In the meta-testing phase, we assessed ten distinct learning rates for each method and reported the best accuracy. To evaluate the precision of rehearsal-based methods, we experimented with batch sizes of 8 and 16 and observed a minimal performance difference depending on the batch sizes. We selected a batch size of 8 as it requires less memory than larger batch sizes.
@@
Evaluation Metrics: Following [4], this study focuses on the accuracy of unseen samples across new categories in CL. We measure the end-to-end latency, energy consumption, and peak memory by deploying various Meta-CL methods using MetaCLBench on edge devices: Jetson Nano, Raspberry Pi 3B+, Raspberry Pi Zero 2. Specifically, we measure system performance metrics by continually learning all given classes for deployed DNNs on embedded devices. Peak memory includes: (1) model memory for updated weights, (2) optimizer memory for gradients and momentum, (3) activations memory for intermediate outputs during weight updates, and (4) rehearsal samples. To measure end-to-end latency and energy consumption, we account for the time and energy required to: (1) load the model and (2) perform Meta-CL using all available samples (e.g., 30) across all classes (e.g., 30 for CIFAR-100 and 10 for GSCv2). Energy consumption is measured by monitoring the power consumption of the edge device using a YOTINO USB power meter. We derive the energy consumption using the equation (Energy = Power x Time).
IV Results & Findings
@@
This section presents comprehensive evaluation results and findings from MetaCLBench.
IV-A Main Results
We now present deployment-oriented findings, focusing on both accuracy and deployment-critical metrics that determine whether methods are viable for IoT deployment. As previous research [4, 12] has shown that ANML often outperforms OML when using 3-layer CNN architectures, we evaluate ANML on CNN architectures. For advanced models like YAMNet and ViT, we prioritized OML evaluation due to their inherent pre-training components, which eliminate the need for additional pre-training through ANML or the AIM module—additions that neither significantly improve accuracy nor justify their computational cost.
Accuracy. Tables I and II present accuracy across all method-dataset combinations. Oracle models (Oracle ANML for CNN, Oracle OML for YAMNet and ViT), trained on all classes simultaneously, establish upper-bound accuracy representing ideal (but unrealistic) deployment scenarios. The key question for IoT practitioners is not which method achieves the highest accuracy, but which methods achieve acceptable accuracy while remaining deployable on target hardware.
By comparing Meta-CL methods against Oracle models, we demonstrate the trade-offs between Meta-CL approaches and ideal scenarios where all data is available initially, highlighting the methods’ efficiency under practical constraints. Also, "Pre-trained" approach (transfer learning applied on pre-trained weights without meta-training) shows lower accuracy, averaging 22.18% and 13.8% across the datasets for Convolutional Neural Networks (CNNs) and advanced architectures (YAMNet and ViT). This result indicates that traditional transfer learning fails to address few-shot learning challenges. In CNN architecture, baseline ANML improved upon the "pre-trained" approach with a 16.7% average accuracy gain. The AIM enhancement further increased accuracy by 2.54%. However, these improvements remained below Oracle and LifeLearner by 15.92%.
Peak Memory Footprint. For IoT deployment, peak memory during learning—not inference—determines viability. Hence, we conducted measurements of the maximum memory requirements for backpropagation and rehearsal samples. The backpropagation memory comprises: (1) model parameters; (2) optimizer states (gradients and momentum vectors); (3) activation (intermediate tensors used during backpropagation).
First, the results in Tables I and II reveal a critical deployment barrier: AIM-enhanced methods (ANML+AIM, OML+AIM) requires 608-2,648 MB (except for a single case), exceeding available RAM on Pi 3B+ and Pi Zero 2, due to additional parameters and activations required during training AIMs. These methods, despite strong accuracy, are non-deployable on typical IoT hardware.
Second, while OML, ANML, Oracle, Pre-trained, and LifeLearner demonstrate lower memory requirements, Pre-trained and ANML are less accurate. LifeLearner emerges as the optimal solution, requiring only 136-496 MB while maintaining accuracy comparable to Oracle. This efficiency is achieved through its compression module for latent replays, reducing activation memory and overall memory footprint, making it particularly suitable for resource-constrained devices.
Deployment Viability. Out-of-memory (OOM) failures occurred for ANML+AIM and OML+AIM on all datasets except GSCv2 and CIFAR-100 when deployed on Pi 3B+. This finding has immediate practical implications: IoT practitioners targeting sub-1 GB devices should exclude AIM-enhanced methods from consideration regardless of their reported accuracy.
@@ IV-A Main Results
TABLE I: Performance and computational costs, memory footprint of six representative Meta-CL methods using CNN on five datasets of image and audio domains. OOM indicates an out-of-memory issue.
@@
CIFIAR-100 Pre-trained 0.260 39.69 MB 305s 1.44 kJ
ANML 0.272 39.69 MB 309s 1.44 kJ
ANML+AIM 0.346 1,093 MB 6,390s 29.43 kJ
OML+AIM 0.311 834.1 MB 1,481s 6.84 kJ
Raw ANML 0.392 99.5 MB 11,424s 52.5 kJ
Oracle ANML 0.445 39.93 MB 1,866s 8.55 kJ
LifeLearner 0.452 15.45 MB 374s 1.71kJ
Mini ImageNet Pre-trained 0.258 474.5 MB 1,198s 5.5 kJ
@@
Oracle ANML 0.438 475.0 MB 3,414s 15.7 kJ
LifeLearner 0.433 137 MB 1,343s 6.17 kJ
GSCv2 Pre-trained 0.213 10.16 MB 71.1s 0.324 kJ
ANML 0.429 10.16 MB 71.1s 0.324 kJ
ANML+AIM 0.710 608.2 MB 394.2s 1.8 kJ
OML+AIM 0.649 135.2 MB 157.5s 0.72 kJ
Raw ANML 0.135 45.72 MB 735.6s 3.37 kJ
Oracle ANML 0.712 10.20 MB 569.7s 2.6 kJ
LifeLearner 0.713 3.40 MB 75.6s 0.35 kJ
Urban Sound8K Pre-trained 0.182 1,382 MB 302s 1.6 kJ
@@
ANML+AIM 0.439 2,593 MB OOM OOM
OML+AIM 0.385 2,648 MB OOM OOM
Raw ANML 0.667 1,456 MB 60,120s 279.0 kJ
@@
ESC-50 Pre-trained 0.196 1,163 MB 1,359s 7.2 kJ
ANML 0.308 1,163 MB 1,274s 7.2 kJ
ANML+AIM 0.233 2,305 MB OOM OOM
OML+AIM 0.181 2,152 MB OOM OOM
Raw ANML 0.358 5,005 MB OOM OOM
Oracle ANML 0.435 1,167 MB 4,120s 20.3 kJ
LifeLearner 0.381 316.3 MB 1,445s 7.8 kJ
TABLE II: Performance and computational costs, memory footprint of six representative Meta-CL methods using ViT and YAMNet on five datasets of image and audio domains.
@@
Pre-trained 0.119 29.8 MB 615s 2.96 kJ
OML 0.271 29.8 MB 628s 3.03 kJ
CIFAR-100 Raw OML 0.40 153 MB 22,968s 103 kJ
Oracle OML 0.361 29.5 MB 3,672s 19.4 kJ
Latent OML 0.365 36.8 MB 764s 3.76 kJ
Mini ImageNet Pre-trained 0.234 337 MB 2,056s 11.2 kJ
OML 0.255 337 MB 2,203s 11.4 kJ
Raw OML 0.454 1,334 MB 301,456s 1150.0 kJ
Oracle OML 0.512 338 MB 6,301s 31.0 kJ
Latent OML 0.376 337 MB 2,250s 13.0 kJ
@@
Pre-trained 0.061 39.2 MB 213.6s 1.08 kJ
OML 0.329 39.2 MB 209.8s 1.08 kJ
GSCv2 Raw OML 0.749 120 MB 1,482s 6.85 kJ
Oracle OML 0.701 39.7 MB 1,685s 14.6 kJ
Latent OML 0.704 40.9 MB 235.8s 0.43 kJ
Urban Sound8K Pre-trained 0.186 39.2 MB 912s 5.4 kJ
OML 0.262 39.2 MB 910s 5.36 kJ
Raw OML 0.595 140 MB 120,965s 550 kJ
Oracle OML 0.729 42.4 MB 2.501s 21.1 kJ
@@
Pre-trained 0.090 39.8 MB 4,052s 19.6 kJ
OML 0.135 39.8 MB 3,986s 9.6 kJ
ESC-50 Raw OML 0.577 210.8 MB 28,465s 39.6 kJ
Oracle OML 0.442 42.5 MB 12,473s 84.2 kJ
Latent OML 0.458 44.3 MB 4,536s 7.4 kJ
End-to-end Latency & Energy. Tables I and II also present the operational efficiency of Meta-CL methods on Raspberry Pi 3B+, focusing on end-to-end latency and energy consumption. The end-to-end latency comprises four components: model loading (5%), inference (30%), backpropagation (61%), and data processing overhead (4%).
@@
First, we measured the end-to-end latency of CL methods on the Raspberry Pi 3B+ across all classes (30 samples per class). LifeLearner achieved fast end-to-end latency. Oracle exhibited higher training latency due to i.i.d. training over multiple epochs until it converges. AIM’s implementation of attention mechanism selection during backpropagation increased computational overhead. LifeLearner reduces training and memory latency through latent replay and model freezing. Due to the large amount of memory required by the AIM series, crashes occurred several times during the test due to memory exhaustion. In particular, Raw ANML requires substantial latency and energy consumption (e.g., 11,424s and 52.5 kJ on CIFAR-100) as it requires full backpropagation of the entire networks.
[…]
Generalizability across IoT Hardware. We extended our evaluation to diverse edge devices with varying computational capabilities (Jetson Nano with enhanced specifications and Pi Zero 2 with reduced capacity). The devices we selected span from entry-level edge AI platform (Jetson Nano: 4GB RAM, quad-core processor) to general IoT device (Pi 3B+: 1GB RAM) to highly constrained (Pi Zero 2: 0.5GB RAM) configurations, covering key points in the resource continuum of IoT platforms. By testing across this hardware spectrum, we observe how different Meta-CL methods perform under varying resource constraints, revealing which approaches are most suitable for specific hardware profiles. Our findings demonstrate that the performance patterns and resource requirements scale predictably with hardware capabilities, allowing practitioners to extrapolate to similar devices. For example, devices with RAM less than 512 MB would likely benefit from the same optimization strategies identified for Pi Zero 2, while those with about 1 GB RAM might follow patterns observed on Pi 3B+. The table III below shows the end-to-end latency and energy consumption results across these three edge devices, confirming that LifeLearner achieves a 7.5-fold reduction in latency compared to Oracle.
(a) UrbanSound8K
(b) ESC-50
@@
TABLE III: End-to-end latency and energy consumption results of running different Meta-CL methods on three different edge devices